A 350 mm diameter grinding wheel must be driven at a grinding speed (speed at circumference) if 1200 meters per minute. It takes 5 seconds to reach this speed from rest, runs at this speed for the next two minutes and then starts to decelerate uniformly, finally coming to rest after a further 30 seconds. Draw a neat angular velocity-time graph and determine how many revolutions the wheel made in total.


Also clearly show what the following values are:


ω1

θ1

θ2

θ3

θtotal

1
Expert's answer
2021-08-18T13:52:02-0400

"\text{diameter} =350mm\\
\text{radius} = 175mm
\\ v_1 = 1200m/min = 20m/s\\
\text{but, } v = \omega r\\
20m/s =\omega_1 × 175 × 10^{-3}m\\
\omega_1 = 114.3\ rad/s"


"\omega_0 = 0\ rad/s
\\ \omega_2 = 0\ rad/s"


For first part of the motion

"\theta_1 = (\dfrac{\omega_1+\omega_0}2 )t"


"\theta_1= \dfrac{114.3+0}2×5= 285.75\text{ rad}"


for second part of the motion

"\theta_2= \omega_1 ×t = 114.3 ×2(60)= 13716\text{ rad}"


for third part of the motion

"\theta_3 =(\dfrac{\omega_2+\omega_1}2 )t"


"\theta_3 = \dfrac{0+114.3}{2} × 30=1714.5 \text{ rad}"


"\text{Total angular revolutions} = 285.75+ 13716+1714.5= 15716.25"


"\text{Number of revolutions } = \dfrac{\theta}{2π}=\dfrac{15716.25}{2π} = 2501.32 \text{ rev}"